(4^x+2) (2^x) = 32
Answer by stanbon(48535)

(4^(x+2)) (2^x) = 32
----------
[2^(2x+4)][2^x] = 2^5
-----------
2^(3x+4) = 2^5
3x+4 = 5
3x + 1
x = 1/3
============
Cheers,
Stan H.